how do i get 2 answers when solving quadratic equations? my teacher said i was going to get two and i need help. -3x-11x-4=0 and please include how you got the answer.
Answer by stanbon(75887)

how do i get 2 answers when solving quadratic equations? my teacher said i was going to get two and i need help. -3x-11x-4=0 and please include how you got the answer.
-------------------------------------
-3x-11x-4=0 is NOT a quadratic equation; it is a linear equation which has
one solution.
-----
-3x-11x-4=0
-14x = 4
x = -2/7
======================
Cheers,
Stan H.